# A PBN-RL-XAI Framework for Discovering a “Hit-and-Run” Therapeutic Strategy in Melanoma

### v4

@@ III Results and Discussions
III-A The Resistant Network is Characterized by a Rigid and Canalized Attractor Landscape
To investigate the dynamic consequences of network rewiring, we performed a computational attractor analysis on both the responder and non-responder PBN models using the optPBN MATLAB toolbox[23]. For each PBN, we estimated the attractor landscape by generating an ensemble of 1,000,000 individual Boolean networks (BNs), with each BN stochastically sampled to represent a specific realization of the probabilistic model. We then identified all steady-state and cyclical attractors for each sampled BN. By aggregating these results across all simulations, we calculated the probability of observing each unique attractor, yielding a quantitative map of the system’s dynamics.
The analysis reveals a reorganization of the network’s dynamic landscape, shifting from a plastic, multi-stable state in responders to a rigid, canalized state in non-responders. The responder network exhibits high plasticity, characterized by 22 distinct attractors, with the most probable one accounting for only 12.55% of the total probability basin. In stark contrast, the non-responder network is highly rigid, collapsing into only 16 attractors. A single singleton attractor dominates this landscape, accounting for 50.02% of the total probability, which indicates the network has been rewired to robustly maintain a specific resistant phenotype.
TABLE I: Key cellular phenotypes identified through attractor analysis. Each state is defined by the binary expression (OFF=0, ON=1) of the core IPRES genes AXL, WNT5A, and ROR2, distinguishing between therapy-sensitive and resistant subtypes.
Phenotype Key Gene States
(AXL, WNT5A, ROR2)
@@ III-A The Resistant Network is Characterized by a Rigid and Canalized Attractor Landscape
Decoding the dominant attractors provides insight into the specific gene expression patterns that define the sensitive and resistant states. While the responder landscape was fragmented across multiple phenotypes, the analysis of the non-responder landscape revealed that the dominant attractor did not fully match the expected Innate anti-PD-1 Resistance (IPRES) signature (e.g., AXL=1, WNT5A=1, ROR2=1). The key identified cellular phenotypes are summarized in Table I.
This analysis demonstrates that innate resistance corresponds to a fundamental system-level transition to a new, highly stable dynamic regime, where the network becomes hard-wired to occupy a dominant, therapy-resistant attractor.
@@ III-B Network Rewiring Establishes a JUN/LOXL2 Regulatory Axis
To identify the key regulatory hubs, we calculated a total influence score for each gene by summing all its incoming and outgoing edge weights from the network’s influence matrix. This score quantifies a gene’s overall impact as both a regulator and a target within the system. The analysis revealed a significant shift in the network’s command structure, establishing the transcription factor JUN as the unequivocal master regulator of the resistant state, with LOXL2 emerging as its primary executive arm. As shown in Figure 1, JUN’s total influence increases dramatically in the non-responder network, and LOXL2 exhibits the second-largest gain. This ascent pinpoints the JUN/LOXL2 axis as the central driver of the switch to the therapy-resistant phenotype.
Figure 1: Shift in total gene influence from the responder (blue) to the non-responder (red) model. The plot highlights the change in influence score for each gene, sorted by their final influence in the non-responder state. The significant increase in JUN’s influence, along with notable shifts in LOXL2, illustrates the systemic rewiring of regulatory control.
@@
III-C Reinforcement Learning Identifies the MAP2K3/LOXL2 Axis as the Optimal Dynamic Control Target
To identify an optimal, unconstrained control policy, we trained a reinforcement learning agent to steer the network from three resistant attractor states to the target sensitive phenotype as in Table I. The resulting policy was exceptionally robust, achieving a 100% success rate over 10,000 evaluation trials.
An analysis of the agent’s actions reveals a highly specific and hierarchical control strategy (Table II). The agent’s policy converged on targeting the MAP2K3/LOXL2 axis; its most frequent intervention was to flip MAP2K3, followed by flipping LOXL2. This demonstrates that while static network analysis pointed to JUN as a master regulator, the dynamic, goal-oriented learning process discovered a more efficient control pathway centered on the upstream kinase MAP2K3.
TABLE II: Action frequencies of the trained reinforcement learning agent over 10,000 evaluation episodes. The data reveals a highly specific control policy that converges on flipping MAP2K3/LOXL2 as the primary intervention, identifying it as the optimal dynamic control target.
Action Times Chosen
Flip Gene ’MAP2K3’ 9037
Flip Gene ’LOXL2’ 3366
Flip Gene ’TAGLN’ 659
Flip Gene ’RELA’ 483
Flip Gene ’WNT5A’ 293
Flip Gene ’JUN’ 183
III-D Temporal Inhibition Reveals a “Hit-and-Run” Therapeutic Strategy
@@
To identify a clinically relevant, transient intervention strategy, we performed a comprehensive analysis simulating a temporary “priming” inhibition of key targets (JUN, LOXL2, and MAPK3) for varying durations. After each priming phase, a pre-trained RL agent was tasked with guiding the network to the therapy-sensitive state.
@@
The results, visualized in Fig. 2, identified a highly effective and specific optimal strategy: a 4-step temporary inhibition of LOXL2, which achieved the highest success rate of 93.45%. The non-monotonic relationship between inhibition duration and success rate for both LOXL2 and JUN provides strong evidence for a “hit-and-run” mechanism, where a precisely timed intervention is more effective than a sustained one. A striking and consistent finding across all experiments was that the agent’s overwhelming post-priming policy was to “Do Nothing”. This validates our hypothesis that a brief, targeted intervention is sufficient to destabilize the resistant network, allowing the system’s own dynamics to complete the transition to a sensitive state.
@@
Figure 2: Efficacy of “hit-and-run” priming strategies. The plot shows the control success rate as a function of the inhibition duration for three key targets. A 4-step inhibition of LOXL2 emerges as the optimal transient strategy, significantly outperforming other interventions.
@@
Figure 3: Comparative SHAP analysis for the agent’s decision to flip MAP2K3 versus JUN from a resistant state. The plot shows the SHAP values for each gene’s contribution to the respective action’s logit, revealing a context-dependent policy.
III-E Mechanistic Policy Explanation via SHAP
To deconstruct the sophisticated policy learned by the reinforcement learning agent, we employed SHAP[22]. This allowed us to move beyond observing the agent’s actions to understanding its underlying reasoning, providing mechanistic explanations for its control strategy.
III-E1 Context-Dependent Intervention Choice
@@
We sought to understand why the agent prioritizes flipping MAP2K3 over other influential genes like JUN. A comparative SHAP analysis reveals a highly context-dependent policy (Fig. 3). While high LOXL2 states push the agent towards flipping either MAP2K3 or JUN, the states of NRAS and TAGLN act as decisive signals: high NRAS encourages MAP2K3 targeting, while high TAGLN favors JUN intervention.
This context-dependent strategy aligns with melanoma signaling biology. NRAS, mutated in 15-20% of melanomas [24], activates the MAPK cascade through MAP2K3 [25]. The agent’s preference for MAP2K3 in NRAS-high contexts reflects understanding of this upstream-downstream relationship, intercepting oncogenic signals at a strategic chokepoint. Conversely, TAGLN drives cytoskeletal remodeling and epithelial-mesenchymal transition [26], prompting the agent to target JUN—a key AP-1 transcription factor [27]—to disrupt transcriptional programs driving resistance.
This demonstrates that the agent discerns specific molecular contexts rather than identifying a generic resistant state, selecting the most effective pathway-specific intervention and mirroring precision oncology strategies [28].
III-E2 Validation of the “Hit-and-Run” Mechanism
Next, we used SHAP to explain why the 4-step inhibition of LOXL2 was more successful than the sub-optimal 2-step inhibition of JUN. We analyzed the network state after each priming strategy by measuring the average SHAP importance of each gene. The results show that the optimal LOXL2 priming is more effective because it does a better job of erasing the “vulnerability signature” that triggers the agent’s desire to intervene (Fig. 4). After the LOXL2 inhibition, the residual SHAP values for key driver genes like NRAS and RELA are significantly lower than after the JUN inhibition, indicating the network has been pushed to a more stable, less-resistant state.
@@ III-E2 Validation of the “Hit-and-Run” Mechanism
Figure 4: Comparison of residual SHAP feature importance after the optimal (4-step LOXL2) versus a sub-optimal (2-step JUN) priming strategy. The more effective LOXL2 priming leaves a state with lower overall SHAP importance, requiring less subsequent intervention.
@@
Figure 5: SHAP trajectory analysis explaining the agent’s policy. The most frequent action was consistently “Do Nothing,” represented by a dash (-) on the x-axis. Persistent negative SHAP values for NRAS and RELA reveal a vulnerability signature that the agent learns to ignore, validating the “hit-and-run” strategy.
III-E3 Dynamic Visualization of the Control Strategy
Finally, to visualize the agent’s reasoning, we analyzed the average of 10,000 simulated 15-step control episodes (Fig. 5). The resulting SHAP trajectory heatmap reveals that the most frequent action at every step was “Do Nothing,” providing strong quantitative support for a “hit-and-run” control mechanism where the network self-corrects after an initial priming.
The heatmap illustrates a persistent “vulnerability signature” that drives the agent’s policy. At the initial step, we observe a powerful negative SHAP contribution from the key driver genes NRAS and RELA, indicating a strong initial impetus for intervention. While this signal weakens after the first step, it does not disappear; instead, it persists at a consistent, lower intensity for the remainder of the episode. This suggests that even as the network evolves, these genes continue to exert a background pressure that the agent has learned to correctly ignore. Other genes, including LOXL2 and MAPK3, contribute weaker, transient negative signals in the early steps before becoming neutral.
This temporal pattern of network self-organization following an initial perturbation aligns with emerging therapeutic paradigms in cancer treatment. The concept parallels successful clinical strategies such as induction chemotherapy in acute leukemias, where intensive initial treatment creates system instability that resolves into remission [34], and immune checkpoint blockade, where brief T-cell activation triggers sustained anti-tumor responses [35]. The identification of NRAS and RELA as key, persistent vulnerability drivers reflects established understanding that RAS/MAPK hyperactivation and NF-$\kappa$B inflammatory signaling represent critical resistance mechanisms [36, 1], supporting the biological plausibility of the discovered intervention strategy.
@@ IV Conclusion
Most significantly, our reinforcement learning approach discovered a non-obvious ”hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving over 93% success in silico. Explainable AI analysis demonstrated that this transient intervention erases the molecular vulnerability signature stabilizing resistance, allowing the network’s intrinsic dynamics to complete the therapeutic transition.
Our findings are based on an in silico model with inherent limitations, including its binary gene representation, focus on a core gene set, and the specific reward configuration for reinforcement learning. For instance, using a positive action cost instead of zero would penalize frequent interventions and could yield a different optimal policy. Consequently, the identified strategies should be interpreted as computational hypotheses that require rigorous experimental validation in cell lines and animal models before any clinical consideration.
The discovery of time-dependent optimal strategies challenges traditional paradigms focused on sustained pathway inhibition. The integrated PBN-RL-XAI pipeline offers a generalizable approach for discovering novel control strategies in complex, treatment-resistant diseases, providing a powerful framework for future therapeutic development.
V Code Availability
All source code developed for the PBN inference, reinforcement learning, and subsequent analyses presented in this paper are publicly available in our GitHub repository at https://github.com/Liu-Zhonglin/pbn-melanoma-project.

### v5

@@ III Results and Discussions
III-A Network Topology Rewiring Drives Attractor Landscape Rigidification
To investigate how network rewiring affects system dynamics, we performed computational attractor analysis on both PBN models using the optPBN MATLAB toolbox[25]. Each PBN was sampled as 1,000,000 individual Boolean networks to map the probability distribution of steady-state and cyclical attractors.
Network topology comparison reveals fundamental structural differences between responder and non-responder networks (detailed network visualizations in Appendix Fig. 6). The non-responder network exhibits higher connectivity (45 vs 41 edges, density 0.341 vs 0.311) with TAGLN functioning as a resistance hub receiving inputs from RELA and JUN while regulating ROR2. Conversely, the responder network shows simplified TAGLN regulation and distinct AXL pathway rewiring toward LOXL2-mediated ECM remodeling.
Quantitative analysis of node-level topology differences (Appendix Fig. 7) identifies nodes with substantial changes: absolute in-degree or out-degree differences $\geq$1, or average mutual information differences $\geq$0.2. TAGLN emerges as the most dramatically altered node, with 3-fold reduced in-degree connectivity in responders (-3 inputs, -1 output), confirming its critical role as a resistance coordination hub that becomes disrupted in treatment-responsive networks.
These topological differences drive a dramatic reorganization of the dynamic landscape. The responder network exhibits high plasticity with 22 distinct attractors, where the most probable accounts for only 12.55% of the total probability basin. In contrast, the non-responder network rigidifies into 16 attractors dominated by a single state capturing 50.02% of probability mass, indicating robust maintenance of resistance through the TAGLN-coordinated regulatory circuit.
TABLE I: Key Cellular Phenotypes Identified Through Attractor Analysis
Phenotype (AXL, WNT5A, ROR2)
@@ III-A Network Topology Rewiring Drives Attractor Landscape Rigidification
This analysis demonstrates that innate resistance represents a system-level transition from plastic, multi-stable dynamics to a rigid, canalized regime where specific regulatory circuits (particularly the TAGLN-ROR2 axis) become hard-wired to maintain resistance. The quantitative topology analysis provides mechanistic insight into how network rewiring patterns determine therapeutic response outcomes.
@@ III-B Network Rewiring Establishes a JUN/LOXL2 Regulatory Axis
To identify the key regulatory hubs, we calculated a total influence score for each gene by summing all its incoming and outgoing edge weights from the network’s influence matrix. This score quantifies a gene’s overall impact as both a regulator and a target within the system. The analysis revealed a significant shift in the network’s command structure, establishing the transcription factor JUN as the unequivocal master regulator of the resistant state, with LOXL2 emerging as its primary executive arm. As shown in Fig. 1, JUN’s total influence increases dramatically in the non-responder network, and LOXL2 exhibits the second-largest gain. This ascent pinpoints the JUN/LOXL2 axis as the central driver of the switch to the therapy-resistant phenotype.
Figure 1: Shift in total gene influence from the responder (blue) to the non-responder (red) model. The plot highlights the change in influence score for each gene, sorted by their final influence in the non-responder state.
@@
III-C Temporal Inhibition Reveals a “Hit-and-Run” Therapeutic Strategy
@@
To identify a clinically relevant transient intervention, we simulated temporary “priming” inhibition of key targets (JUN, LOXL2, MAPK3) for varying durations, followed by RL-guided control to the therapy-sensitive state.
@@
The results (Fig. 2) identified an optimal strategy: 4-step LOXL2 inhibition achieving 93.45% success rate. The non-monotonic relationship between inhibition duration and success for both LOXL2 and JUN demonstrates a “hit-and-run” mechanism where precisely timed intervention outperforms sustained targeting. Critically, the agent’s overwhelming post-priming policy was “Do Nothing,” validating that brief intervention destabilizes resistance, allowing natural dynamics to complete transition.
@@
Figure 2: Efficacy of “hit-and-run” priming strategies. A 4-step inhibition of LOXL2 emerges as the optimal transient strategy, significantly outperforming other interventions.
III-D Mechanistic Policy Explanation via SHAP
We employed SHAP analysis[24] to understand the agent’s control reasoning, providing mechanistic explanations for its strategy.
@@ III-D Mechanistic Policy Explanation via SHAP
Figure 3: Comparative SHAP analysis for the agent’s decision to flip MAP2K3 versus JUN from a resistant state.
III-D1 Context-Dependent Intervention Choice
@@
SHAP analysis reveals why the agent prioritizes MAP2K3 over JUN through context-dependent logic (Fig. 3). While high LOXL2 states favor flipping either MAP2K3 or JUN, NRAS and TAGLN states determine the choice: high NRAS directs MAP2K3 targeting, while high TAGLN favors JUN intervention.
This strategy aligns with melanoma biology. In NRAS-mutant contexts (15-20% of melanomas [26]), the agent targets MAP2K3 to intercept MAPK cascade activation [27] at a strategic chokepoint. Conversely, when TAGLN drives epithelial-mesenchymal transition [28], the agent targets JUN [29] to disrupt resistance-associated transcriptional programs. This demonstrates pathway-specific intervention selection mirroring precision oncology strategies [30].
III-D2 Validation of the “Hit-and-Run” Mechanism
SHAP analysis explains why 4-step LOXL2 inhibition outperformed 2-step JUN inhibition by measuring residual “vulnerability signatures” (Fig. 4). LOXL2 priming more effectively erases intervention triggers, leaving significantly lower SHAP values for key drivers like NRAS and RELA, indicating the network reached a more stable, less-resistant state.
@@ III-D2 Validation of the “Hit-and-Run” Mechanism
Figure 4: Residual SHAP importance after optimal (4-step LOXL2) versus sub-optimal (2-step JUN) priming. Effective LOXL2 priming requires less subsequent intervention.
LOXL2’s superior efficacy reflects its pleiotropic role beyond matrix crosslinking—it acts as transcriptional regulator promoting epithelial-mesenchymal transition and coordinates multiple resistance pathways including hypoxia response and immune evasion[31]. This aligns with its position as a master coordinator of resistance-associated programs[32].
III-D3 Dynamic Visualization of the Control Strategy
Analysis of 10,000 simulated episodes reveals persistent “vulnerability signatures” driving the agent’s policy (Fig. 5). The most frequent action was consistently “Do Nothing,” providing quantitative support for “hit-and-run” control where networks self-correct after initial priming.
The heatmap shows powerful initial negative SHAP contributions from NRAS and RELA (-0.4 to -0.6), indicating strong intervention impetus. This signal persists at lower intensity (-0.2 to -0.4) throughout episodes, suggesting these genes exert background pressure the agent learns to ignore. Other genes contribute weaker, transient signals before becoming neutral.
Figure 5: SHAP trajectory analysis showing persistent vulnerability signatures. “Do Nothing” was most frequent action, with NRAS/RELA negative signals that agent learns to ignore, validating the “hit-and-run” strategy.
@@ IV Conclusion
Most significantly, our reinforcement learning approach discovered a non-obvious “hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving over 93% success in silico. Explainable AI analysis demonstrated that this transient intervention erases the molecular vulnerability signature stabilizing resistance, allowing the network’s intrinsic dynamics to complete the therapeutic transition.
Our findings have important limitations. The binary gene representation oversimplifies continuous expression dynamics and misses intermediate states. Our 12-gene focus, while IPRES-motivated, excludes tissue-specific factors and broader regulatory networks. The PBN framework’s discrete time steps and Boolean logic may not capture biological stochasticity and temporal dynamics. The model lacks spatial considerations, metabolic constraints, and immune interactions critical for in vivo efficacy. Additionally, zero action costs in our RL formulation could yield different policies under alternative reward structures. This framework has only been tested on a single melanoma dataset, requiring validation across diverse cancer types and genetic backgrounds.
Clinical translation requires systematic validation: experimental testing in cell lines and organoids to confirm LOXL2 timing effects, preclinical studies in xenograft models combining temporal inhibition with immunotherapy, and development of biomarkers for patient selection. Clinical trials must establish optimal timing protocols, design adaptive frameworks for schedule optimization, and carefully monitor safety and resistance emergence. Phase I studies should focus on pharmacodynamics and immune activation patterns.
The discovery of time-dependent optimal strategies challenges traditional sustained inhibition paradigms. While requiring extensive validation, the integrated PBN-RL-XAI pipeline offers a generalizable computational approach for discovering novel control strategies in complex, treatment-resistant diseases, providing a powerful framework for future therapeutic development.